OpenStudy (anonymous):

f(x) = x + (128) / (x^2+16). Identify all relative extrema. Can anyone help me?

OpenStudy (jennychan12):

r u in calculus?

OpenStudy (anonymous):

yes?

OpenStudy (jennychan12):

okay. sorry find f'(x) then set f'(x) = 0

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Ok So it would be ((254x) / (x^2 + 16)^2) = 0

OpenStudy (anonymous):

This is the point where I originally got, but what do I do once it equals zero?

OpenStudy (jennychan12):

isn't f'(x) = 1-128/(x^2+16)^2 ?

OpenStudy (anonymous):

I am not sure.. I got ((1 - 254x) / (x^2 + 16)^2) = 0

OpenStudy (jennychan12):

how? can u show ur work?

OpenStudy (anonymous):

Could you tell me first what I would do once I make it equal to 0?

OpenStudy (jennychan12):

wait. ur f'(x) is wrong.

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

Try it this way by rewriting, then finding derivative \[x+128(x ^{2}+16)^{-1}\]

OpenStudy (anonymous):

Alright well 1-128/(x^2+16)^2 .. What do I do next?

OpenStudy (anonymous):

Let\[x ^{2}+16\]be u, find derivative of u

OpenStudy (anonymous):

Alright guys, I know how to derive it, the problem is actually finding the relative extrema. So I let the derivative equal zero..then what?

OpenStudy (jennychan12):

1-128/(x^2+16)^2 = 0 u'll have to do some algebra here (common denominator...)

OpenStudy (anonymous):

The problem is, we don't agree that you have correct derivative of the thingie.

OpenStudy (anonymous):

well if I knew what I do when it equals zero I can figure out if it's really right or wrong.

OpenStudy (anonymous):

You have to solve for x.

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

I redid it and I got 1 - 128(2x) / (x^2 + 16)^2

OpenStudy (anonymous):

In the answer choices it says that it is either 1.182, and 4 or -1.182, and -4

OpenStudy (jennychan12):

first u need to find common denominator before solving for 1 - 128(2x) / (x^2 + 16)^2 = 0

OpenStudy (anonymous):

not really sure about this one, would it be (x^2 + 16)^2?

OpenStudy (jennychan12):

yeah. but the 2x shouldn't be there.

OpenStudy (anonymous):

oh I see, so when everything equals to 0, what exactly am I supposed to be doing? This part always confuses me.

OpenStudy (jennychan12):

it should just be 1 - 128/ (x^2 + 16)^2 = 0

OpenStudy (anonymous):

It's been a while since I did derivatives so I punched it in Wolfram \[1-\frac{ 256x }{ (x ^{2}+16)^{2} }\]

OpenStudy (anonymous):

That's what I was trying to say the derivative was^

OpenStudy (anonymous):

Alright So what would I do when it equals 0.. Can you give me the results/process of how to do it.

OpenStudy (jennychan12):

(x^2+16)^2-128 = 0 we don't really care about the denominator because if the numerator equals 0 then the equation equals 0

OpenStudy (anonymous):

Shouldn't 128 by 256x?

OpenStudy (anonymous):

Shouldn't it 'be'

OpenStudy (anonymous):

yea I was saying that earlier.. because of when you use the quotient rule.

OpenStudy (jennychan12):

oh wait sorry yeah it should. i was thinking something else. okay so f'(x) = 1-256x/(x^2+16)^2 so then set it = 0 (x^2+16)^2-256x = 0 and solve.

OpenStudy (anonymous):

Hmm..How would I go about solving this? Since I am solving for x do I have to expand (x^2 + 16)^2?

OpenStudy (anonymous):

Multiply each term by (x^2 + 16)^2 to remove fraction

OpenStudy (jennychan12):

yeah you have to expand (x^2+16)^2

OpenStudy (anonymous):

Alright so I have x^4 + 16x^4 + 256 -256x = 0 at the moment

OpenStudy (anonymous):

any idea what I should focus on next?

OpenStudy (anonymous):

chag?

OpenStudy (anonymous):

My algebra might be wrong, I got x=0, -32, -256. If it is, or if it were right you would do something like this